OpenStudy (anonymous):

2x + 9√(x) - 5 = 0 ; need help solving for x

OpenStudy (amistre64):

its a quadratic; disguised

OpenStudy (amistre64):

[√(x)]^2 = x

OpenStudy (anonymous):

would i have to sqare everything else too?

OpenStudy (amistre64):

no, its just the "form" that you are interested in; the structure of the variable is unimportant at the moment; solve it for: √(x) = quadratic formula outputs

OpenStudy (amistre64):

then determine x from that

OpenStudy (anonymous):

alright thanks

OpenStudy (amistre64):

if its right in me head i see it as: 2(5) = 10 10-1 = 9 (√(x)+10/2)(√(x)-1/2) (√(x)+5)(√(x)-1/2) √(x) = -5 or 1/2 [√(x) = -5 or 1/2] ^2 x = 25 or 1/4

OpenStudy (amistre64):

the 25 might be a bad result that comes with the territory :)